Solve the equation in the interval [0°, 360°).
sin^2θ - sin θ - 12 = 0

sin 2θ = -sin θ

2 cos2θ + 7 sin θ = 5
Answer #1

-4) =0 1 Sin2o-sino-12=0 Sino +3 Sino-usino-12=0 Sien o (Sino+3)-4 (Sino +3)=0 (seno-u) (sin 0+3) = 0 (Sino or (sino +3) =o Sco 2 cos 20 +7. sino =s 211-2 Sino) + 7 Simo = 5 2-4 sin2o + 7 sino =s -usin²o+ 7 sino = 3 4 sin²o -7 sino +3=0 4 sin²o-Brino